OpenStudy (anonymous):

How many times is the mass of one aluminium atom is heavier than one oxygen atom? [ Ar: Al=27, O=16]

OpenStudy (mani_jha):

The atomic mass of Al is 27. The atomic mass of Oxygen is 16. So, what is the ratio of their weights?

OpenStudy (anonymous):

erkk.. don't know ---blur--

OpenStudy (mani_jha):

\[W(Al)/W(O)=27/16\] \[W(Al)=(27/16)W(O)\] So, Aluminium is 27/16 times as heavy as O. Did you understand?

OpenStudy (anonymous):

oh... ok~ i'm understand now